OpenStudy (anonymous):

What is the slope between the points (3, -2) and (8, 4)?

OpenStudy (anonymous):

slope = y2-y1/x2-x1 (3, -2) and (8, 4) x1 y1 x2 y2 plug it in.. what do u get?

OpenStudy (anonymous):

(3,-2) (X1, y1) (8,4) (X2,y2) M=y2-y1/x2-x1 M=4-(-2)/8-3 M=4+2/8-3 M=6/5

OpenStudy (anonymous):

By the way m is slope

OpenStudy (anonymous):

i got -5/6 ... thats why i was confused thanks

OpenStudy (anonymous):

Well now ur all clear what do u think of my answer?

OpenStudy (anonymous):

 so slope=6/5 :)